The number of days a student is absent in a four week period is recorded along with the test score over the material covered over those four weeks. Consider the output from Excel of a linear regression:
Claim: There is a linear correlation between the number of days missed and the test score over the material covered.

Regression Statistics

Multiple R .699389593
R squared .489145802
Adjusted R Square .446574619
Standard error 14.49517418
error observations 14

Anova

df ss ms f
Regression 1 2414.179 2414.179 11.49007
residual 12 2521.321 210.1101
total 13 4935.5
Coefficients Standard Error t Stat P-value
Intercept Classes 82.30589096 5.209439 15.79938 2.14E-09
missed -3.51664837 1.03751 -3.3897 .005371

What is the correlation coefficient, p value, and degrees of freedom?
Answer #1
  • The correlation coefficient is given by 0.6994
  • Here we have to test the claim that There is a linear correlation between the number of days missed and the test score over the material covered.
  • HYPOTHESIS TO BE TESTED.
  • Let the null hypothesis be that there is no linear correlation between the number of days missed and the test score over the material covered.
  • Let the alternate hypothesis be that there is a linear correlation between the number of days missed and the test score over the material covered.
  • Ho:ho =0 vs H1:ho eq 0
  • TEST STATISTICS
  • V2 t2,a) t(n-2
  • 0.6994V14-2 1- 0.6994=3.3897
  • PVALUE
  • By using r code p value is given by
  • Since this is a two sided test
  • p.value = 2*pt(t.value, df=length(lengths-2),lower=FALSE)
  • 2*pt(3.3897, 12,lower=FALSE)
  • =0.005370981
  • CONCLUSION
  • The p value calculated is less than 5% significant level.
  • Hence we do not have enough evidence to reject the null hypothesis.
  • Hence we conclude that  there is a linear correlation between the number of days missed and the test score over the material covered.